# Identifying Sparse Treatment Effects

### v1

@@ 1 Introduction
In this paper, we consider the problem of causal inference from experimental data with high-dimensional outcomes. Our emphasis is on the development of a practical technique that allows the scientist to consider the entire high-dimensional outcome, and yet successfully identify and estimate sparse treatment effects. We present a practical, easily deployed method that identifies the subset of outcome representations that contain the treatment effect given a large pool of outcome representations, and then carries out inference for the treatment effect on the selected subset.
In the remainder of this section, we present some motivation for our approach with examples from real-world experiments and semi-synthetic experiment. First, we observe that for interpretability, instead of working with the raw outcomes, scientists often study a compressed representation of the raw data. A common practice is to compress the outcomes into a one-dimensional representation of interest, $Y\in\mathbb{R}$; alternatively, one can consider $p$ one-dimensional representations, concatenated as $Y\in\mathbb{R}^{p}$, where $p$ is small, and employ multiple testing. This practice is illustrated in the examples below.
@@
    Time series data from wearable devices. Over the last decade, continuous glucose monitoring (CGM) has become an essential component of diabetes management for many people with type 1 diabetes. The study by Aleppo et al. [2017] investigates whether the use of CGM without confirmatory blood glucose monitoring (BGM) measurements is as safe and effective as using CGM adjunctive to BGM in adults with well-controlled type 1 diabetes through a randomized controlled experiment. Their primary outcome was time-in-range, abbreviated as TIR, (the percentage of time when the patient’s blood glucose level lies between 70–180 mg/dL) over the 26-week period of monitoring. In this study, the raw outcome $Y^{\text{raw}}$ is a time series of glucose measurement over 26 weeks, and their choice of one-dimensional representation of outcomes $Y_{1}$ is time-in-range (70–180 mg/dL) over the 26-week period.
@@
    Neuroimaging data. Neuroscientists often investigate the impact of external stimuli on brain activity using neuroimaging techniques, such as functional magnetic resonance imaging (fMRI). In such cases, outcomes are high-dimensional in the form of images or time series of images.
    For example, Koenig et al. [2017] studies the effects of the insulin sensitizer metformin on Alzheimer’s disease through a randomized placebo-controlled crossover design. Arterial Spin Label MRI data, which measures cerebral blood flow (CBF) in different brain regions, was collected at week 0 as baseline and at week 8. Then the researchers tested whether the treatment had significant effects on CBF in 21 pre-defined regions of interest. In this study, the raw outcome $Y^{\text{raw}}$ is a pre-processed brain image and their choice of representation of outcomes, $Y\in\mathbb{R}^{21}$, is measured CBF in 21 regions of interest.
    In another example, Anand et al. [2005] studies the effects of the antidepressant sertraline on corticolimbic connectivity using a matched pairs experimental design. They obtained a series of fMRI scans (512 time points) at week 0 as baseline and week 6. Then the researchers compared correlation coefficients between the time series of four different regions of interest and tested whether there are significant differences between the control and treatment groups. In this study, the raw outcome $Y^{\text{raw}}$ is a time series of brain images and their choice of representation, $Y\in\mathbb{R}^{\binom{4}{2}=6}$, is a vector of correlation coefficients between four different regions of interest.
This practice of considering one or a few outcome representations is common when scientists are interested in a few specific representations based on their scientific motivation. However, if scientists are interested in unraveling how the treatment effect may manifest in the high-dimensional outcome space and picking the right representation, this practice may fail to capture the treatment effect when the treatment effect is sparse.
@@ 1.1 Our Objective and Approach
We consider a setting where the treatment effect is sparse, meaning that the treatment effect manifests in only a few directions in the high-dimensional outcome space $\mathbb{R^{d}}$. Moreover, we suppose that the scientist does not know which outcome representation can effectively capture the treatment effect. In the example in the previous section, this is when a scientist is uncertain about both the duration and the location of the treatment effect.
Without knowing relevant representations, we suppose that scientists consider a vector of $p$ different representations, where $p$ is large. If $p$ is small, it is more likely that a vector of representations does not include representations that can capture the sparse treatment effect. However, if $p$ is large, then even though there is a higher chance of inclusion of representations that capture the treatment effect, the multiple testing correction factor can be substantial.
Our main contribution in this paper is a practical method for identification of the representations where a treatment effect manifests when the number of representations $p$ is large. In our setting, we assume that scientists have $p$ one-dimensional outcome summaries (i.e., representations) $Y\in\mathbb{R}^{p}$ where $p$ can be even larger than $n$. This large pool of outcome summaries may be defined using background knowledge. There are also several representation learning methods for high-dimensional data including time-series and image data sets [Bengio et al., 2013].
@@
Because $p$ is large, we consider a sample splitting method. In the first split, the data scientist finds a large pool of outcome summaries and selects a subset in which the treatment effect may manifest. In the second split, they estimate the treatment effect on the selected subset. Our method primarily focuses on the subset selection process: how to choose a subset of outcome representations in which the treatment effect may manifest. We introduce a sparse regression approach for subset selection, where we employ a regularized weighted linear regression with carefully chosen weights. We show that our method can recover the subset of outcome representations consistently. From simulated experiments, we demonstrate the effectiveness of our method, particularly when the treatment effect is relatively weak.
Before continuing to the technical presentation, below we revisit the examples from real-world experiments to illustrate how this subset selection process can be applied.
  1. 1.
    Time series data from wearable devices (continued). In diabetes treatment, certain interventions may be effective within specific time ranges, such as mornings, post-meal periods, or during sleep. Alternatively, treatments might be effective under specific hypoglycemia or hyperglycemia conditions. In such scenarios, the treatment effect may be reflected in time-in-range (TIR) metrics defined within particular glucose ranges and time frames. If a data scientist is uncertain about the precise glucose levels and time periods affected by the treatment, they may consider various TIR metrics targeting different glucose ranges and time frames, resulting in a vector of outcome representations, $Y\in\mathbb{R}^{p}$, where $p$ is large. If the treatment effect is sparse, using subset selection methods, one may identify a subset of outcome representations where the treatment effect potentially manifests.
  2. 2.
    Neuroimaging data (continued). If scientists aim to gain a more comprehensive understanding of the impact of drug treatment on the brain, they might explore a larger number $p$ of brain regions of interest. For large $p$, the dimension of outcome representations becomes $p$ and $p(p-1)/2$ in Koenig et al. [2017] and Anand et al. [2005], respectively. If we set aside the sample size issues mentioned in the original papers and assume we have sufficient sample sizes, it is possible, once again, to identify a subset of outcome representations where the treatment effect potentially manifests utilizing subset selection methods.
@@ 1.2 Outline of the Paper
In Section 2, we discuss related work. In Section 3, we introduce the formal setting of the paper. In Section 4, we provide an overview of a baseline approach for how a data scientist might select a subset of outcomes where the treatment effect potentially manifests. Section 5 outlines our proposed method based on sparse regressions. Section 5.1 provides the insight behind the proposed procedure, while Section 5.2 presents theoretical results. We conclude in Section 6 where we evaluate the performance of our proposed procedure on synthetic and semi-synthetic data sets.
@@ 2 Related Work
This work falls into the general randomized controlled setting in causal inference [Imbens and Rubin, 2015]. In randomized controlled experiments, linear regression has been employed to improve the precision of average treatment effect estimators through linear regression adjustments with pre-treatment covariates [Deng et al., 2013, Lin, 2013, Jin and Ba, 2023]. Bloniarz et al. [2016] and Wager et al. [2018] have a discussion of high-dimensional regression adjustments when there are a large number of pre-treatment covariates. Our paper also studies high-dimensional regression techniques within randomized controlled experiments, but we focus on the setting where outcomes are high-dimensional instead of pre-treatment covariates.
One of the most popular high-dimensional linear regression technique is the Lasso introduced by Tibshirani [1996]. Conditions for oracle inequalities for the Lasso have been studied by Bunea et al. [2007], van de Geer [2008], Zhang and Huang [2008], Meinshausen and Yu [2009], and Bickel et al. [2009]. Furthermore, variable selection properties of the Lasso have been investigated by Meinshausen and Bühlmann [2006], Zhao and Yu [2006], Lounici [2008], and Wainwright [2009]. In addition, sample-splitting and subsequent statistical inference procedures have been developed in Wasserman and Roeder [2009] and Meinshausen et al. [2008].
@@ 3 Preliminaries
We consider a setting with $n$ individuals. Each individual receives a single binary treatment of interest $T_{i}\in\{0,1\}$, and we observe a subsequent outcome vector, denoted by $Y_{i}\in\mathbb{R}^{p}$, for each individual. (Note that in our motivating examples, this $p$-dimensional outcome is a vector of different representations derived from a raw observed outcome vector $Y_{i}^{\text{raw}}$ $\in\mathbb{R}^{d}$; in our development, we treat $Y_{i}$ as our primary outcome.) In our setting, the outcome representation is high-dimensional: we allow the dimension $p$ to increase with $n$. In our subsequent presentation, we will also consider the inclusion of pre-treatment covariates $X_{i}\in\mathbb{R}^{m}$ for each individual.
@@
Here, we review the standard approach to estimating and inferring $\tau$ when the dimension of the outcome, $p$ , does not increase at the same rate as $n$ and $p\ll n$. The standard approach to estimation of $\tau$ is to consider the difference-in-means estimator:
@@
Inference can be made for each component $j$ of the outcome, $Y_{j}$, with the usual z-test and multiple testing. The p-value for testing the null hypothesis $H_{0j}:\tau_{j}=0$ that there is no treatment effect on $Y_{j}$, is defined as
$P\left(Z\geq\sqrt{n}(\hat{\Sigma}_{\text{DiM},jj})^{-1/2}|{\hat{\tau}_{\text{DiM}}}_{j}|\right)\cdot p,$
where $Z$ is the standard Gaussian random variable and $p$ is a multiple testing correction factor.
Inference can also be made for a group of outcomes. This involves the following statistic, referred to as the Hotelling $T^{2}$ statistic, which generalizes the usual $t$-test:
@@
Under the null hypothesis $H_{0}:\tau=0$ that there is no treatment effect, the Hotelling $T^{2}$ statistic in (5) follows the chi squared distribution with $p$ degrees of freedom. Thus, the p-value for testing the null hypothesis $H_{0}$ is defined as
@@
where $\chi^{2}(p)$ is the chi-squared random variable with $p$ degrees of freedom.
@@
We conclude this section by noting that, as discussed in the introduction, we are primarily interested in settings where the treatment effect may be sparse within the $p$-dimensional outcome. Formally, throughout the remainder of the paper, we suppose that there exists a set of dimensions $S_{\tau}$ such that the sparsity index of $S_{\tau}$ is $|S_{\tau}|:=s_{\tau}<p$ and
$\mathbb{E}[Y_{j}|T=1,X]=\mathbb{E}[Y_{j}|T=0,X]\quad$ for $\text{for }j\notin S_{\tau},$
$\mathbb{E}[Y_{j}|T=1,X]\neq\mathbb{E}[Y_{j}|T=0,X]\quad$ for $\text{for }j\in S_{\tau}.$
Our goal in this paper is to present a practical regression-based method for approximate recovery of $S_{\tau}$, together with valid inference on treatment effects localized to the identified subset. In Section 1, we provided motivation for the use of a sample-splitting approach, where half the data sample is used to identify $S_{\tau}$, and the second half of the data sample is used for inference. Next, we present a simple, commonly employed baseline approach for this purpose (first without, then with pre-treatment covariates); we subsequently compare our proposed approach to this baseline.
@@ 4 A Baseline Approach
In this section we present a simple, commonly used baseline approach that serves as a reference point. For simplicity, we assume that the data scientist knows (or has a bound) on the sparsity $s_{\tau}$.
The most common method employed for subset identification is to rank outcomes by their treatment effect sizes, and select outcomes with the largest effect sizes. This is equivalent to considering the following objective:
$\hat{\beta}=\operatorname*{argmin}_{\beta}\Bigg{|}\Bigg{|}\left(\frac{{\hat{\tau}_{\text{DiM}}}_{j}}{{\sqrt{\hat{\Sigma}_{\text{DiM}}}_{jj}}}\right)_{j=1}^{p}-\beta\text{ }\Bigg{|}\Bigg{|}_{2}^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau},$ (7)
where $||\beta||_{0}=\sum_{j=1}^{p}1\{\beta_{j}\neq 0\}$ is the $\ell_{0}$ norm of $\beta$. Then, the estimated $\hat{S}_{\tau}$ is defined as
@@
which is simply the top $s_{\tau}$ outcomes based on the effect sizes of average treatment effects.
Having identified a subset $\hat{S}_{\tau}$ on the first data split, inference can proceed as in Section 3 on the second data split, but restricted to the set $\hat{S}_{\tau}$. We can specialize the definitions in (1) and (4) to the subset $\hat{S}_{\tau}$, and carry out inference using the second data split as suggested in (3) or (6) defined with these $\hat{S}_{\tau}$-specific quantities. Note that in (3), we would still apply a correction factor for multiple testing, but it is significantly smaller, especially when $|\hat{S}_{\tau}|$ is much smaller than $p$.
Going beyond the basic method above, in causal inference, the pre-treatment covariates are often leveraged to reduce the variance of the treatment effect estimators (Deng et al. [2013], Lin [2013], Jin and Ba [2023]). As an example, the CUPED estimator [Deng et al., 2013] estimates $\tau$ as

### v2

@@ 1 Introduction
In this paper, we consider the problem of causal inference from experimental data with high-dimensional outcomes. Our emphasis is on the development of a practical technique that allows the scientist to consider high-dimensional outcome summaries, and yet successfully identify and estimate sparse treatment effects. We present a practical, easily deployed method that identifies the subset of outcome summaries that captures the sparse treatment effect given a large pool of outcome summaries, and then carries out inference for the treatment effect on the selected subset.
In the remainder of this section, we present some motivation for our approach with examples from real-world and semi-synthetic experiments. First, we observe that for interpretability, instead of working with the raw high-dimensional outcomes, scientists often study a compressed summary of the raw data. A common practice is to compress the outcomes into a one-dimensional summary of interest, $Y\in\mathbb{R}$; alternatively, one can consider $p$ one-dimensional summaries, concatenated as $Y\in\mathbb{R}^{p}$, where $p$ is small, and employ multiple testing. This practice is illustrated in the examples below.
@@
    Time series data from wearable devices. Over the last decade, continuous glucose monitoring (CGM) has become an essential component of diabetes management for many people with type 1 diabetes. The study by Aleppo et al. [2017] investigates whether the use of CGM without confirmatory blood glucose monitoring (BGM) measurements is as safe and effective as using CGM adjunctive to BGM in adults with well-controlled type 1 diabetes through a randomized controlled experiment. Their primary outcome was time-in-range (the percentage of time when the patient’s blood glucose level lies between 70–180 mg/dL) over the 26-week period of monitoring. In this study, the raw outcome is a time series of glucose measurement over 26 weeks, and their choice of one-dimensional outcome summary $Y$ is time-in-range over the 26-week period.
@@
    Neuroimaging data. Neuroscientists often investigate the impact of external stimuli on brain activity using neuroimaging techniques, such as functional magnetic resonance imaging (fMRI). In such cases, outcomes are high-dimensional in the form of images or a time series of images.
    For example, Koenig et al. [2017] studied the effects of the insulin sensitizer metformin on Alzheimer’s disease through a randomized placebo-controlled crossover design. Arterial Spin Label MRI data, which measures cerebral blood flow (CBF) in different brain regions, was collected at week 0 as baseline and at week 8. Then the researchers tested whether the treatment had significant effects on CBF in 21 pre-defined regions of interest. In this study, the raw outcome is a pre-processed brain image and their choice of outcome summaries, $Y\in\mathbb{R}^{21}$, is measured CBF in 21 regions of interest.
    In another example, Anand et al. [2005] studied the effects of the antidepressant sertraline on corticolimbic connectivity using a matched pairs experimental design. They obtained a series of fMRI scans (512 time points) at week 0 as baseline and week 6. Then the researchers compared correlation coefficients between the time series of four different regions of interest and tested whether there are significant differences between the control and treatment groups. In this study, the raw outcome is a time series of brain images and their choice of outcome summaries, $Y\in\mathbb{R}^{6}$, is a vector of correlation coefficients between four different regions of interest.
This practice is common when scientists are interested in a few specific outcome summaries based on their scientific motivation. However, if scientists are interested in unraveling how the treatment effect may manifest in the high-dimensional outcome space, this practice may fail to capture the treatment effect when the treatment effect is sparse.
@@
The preceding discussion motivates the main challenges we address in this paper. First, given a large number of possible outcome summaries (e.g., windows of different start time and duration), where is a treatment effect most likely present? Second, what is an efficient means of leveraging pre-treatment covariates (e.g., the prior week’s CGM time series) to improve detection of appropriate outcome summaries? And third, what is the performance of such an approach, both in terms of recovery of the appropriate outcome summaries, as well as in terms of statistical power?
@@ 1.1 Our Objective and Approach
We consider a setting where the outcome of interest lies in the high-dimensional space $\mathbb{R}^{p}$, with $p$ possibly greater than $n$. As in the examples in the previous section, the outcome of interest may be $p$ one-dimensional summaries of the raw outcome. We assume that the treatment effect is sparse, meaning that the treatment effect manifests in only a few directions in the high-dimensional space $\mathbb{R}^{p}$. Additionally, the scientist does not know which dimensions effectively capture the treatment effect.
Our main contribution in this paper is a practical method for identification of the subset of outcome summaries in which a treatment effect manifests when the dimension $p$ is large. We propose a two-step procedure given two data splits. For example, in a clinical setting, we often begin with a small-scale experiment as pilot or proof-of-concept (POC) study, followed by a much larger clinical trial. Alternatively, these two data splits can also be obtained through sample-splitting. In the first stage, the data scientist defines a large pool of outcome summaries and selects a subset where the treatment effect may manifest using the first data split. In the second stage, the second data split is used to estimate the treatment effect on this selected subset.
@@
Our method primarily focuses on the subset selection process: how to choose a subset of outcome summaries in which the treatment effect may manifest. We introduce a sparse regression approach for subset selection, where we employ a regularized weighted linear regression with carefully chosen weights. We show that our method can recover the subset of outcome summaries consistently. From simulated experiments, we demonstrate the effectiveness of our method, particularly when the treatment effect is relatively weak and where $p\gg n$ during the subset selection phase. In our clinical applications of interest, we typically face $p\gg n$ during the initial POC study due to the small number of subjects. In other settings, such as analysis of gene expression data, the massively high-dimensional outcomes create a $p\gg n$ scenario even when $n$ is still large enough to perform sample-splitting.
@@ 1.2 Outline of the Paper
In Section 2, we discuss related work. In Section 3, we introduce the formal setting of the paper. In Section 4, we provide an overview of a baseline approach for how a data scientist might select a subset of outcome summaries where the treatment effect potentially manifests. Section 5 outlines our proposed method based on sparse regressions. Section 5.1 provides technical insights behind the proposed method and Section 5.2 presents theoretical results. We conclude in Section 6 where we evaluate the performance of our proposed procedure on synthetic and semi-synthetic data sets.
@@ 2 Related Work
This work considers the general randomized controlled setting in causal inference [Imbens and Rubin, 2015]. In randomized controlled experiments, linear regression has been employed to improve the precision of average treatment effect estimators through linear regression adjustments with pre-treatment covariates [Deng et al., 2013, Lin, 2013, Jin and Ba, 2023]. Bloniarz et al. [2016] and Wager et al. [2018] have a discussion of high-dimensional regression adjustments when there are a large number of pre-treatment covariates. Our paper also studies high-dimensional regression techniques within randomized controlled experiments, but we focus on the setting where outcomes are high-dimensional instead of pre-treatment covariates.
One of the most popular high-dimensional linear regression techniques is the Lasso introduced by Tibshirani [1996]. Conditions for oracle inequalities for the Lasso have been studied by Bunea et al. [2007], van de Geer [2008], Zhang and Huang [2008], Meinshausen and Yu [2009], and Bickel et al. [2009]. Furthermore, variable selection properties of the Lasso have been investigated by Meinshausen and Bühlmann [2006], Zhao and Yu [2006], Lounici [2008], and Wainwright [2009]. In addition, sample-splitting and subsequent statistical inference procedures have been developed in Wasserman and Roeder [2009] and Meinshausen et al. [2008].
@@ 3 Preliminaries
We consider a setting with $n$ individuals. Each individual receives a single binary treatment of interest $T_{i}\in\{0,1\}$, and we observe a subsequent vector of outcome summaries, denoted by $Y_{i}\in\mathbb{R}^{p}$, for each individual. In our setting, the primary outcome, a vector of outcome summaries, is high-dimensional: we allow the dimension $p$ to increase with $n$. In our subsequent presentation, we will also consider the inclusion of pre-treatment covariates $X_{i}\in\mathbb{R}^{m}$ for each individual.
@@
Here, we review the standard approach to estimating $\tau$ and quantifying uncertainty when the dimension of the outcome, $p$, does not increase at the same rate as $n$ and $p\ll n$. The standard approach to estimation of $\tau$ is to consider the difference-in-means estimator:
@@
Inference can be made for each component $j$ of the outcome, $Y_{.,j}$, with the usual z-test and multiple testing. The p-value for testing the null hypothesis $H_{0j}:\tau_{j}=0$ that there is no treatment effect on $Y_{.,j}$, is defined as
$P\left(Z\geq\sqrt{n}(\hat{\Sigma}_{\text{DiM},jj})^{-1/2}|{\hat{\tau}_{\text{DiM}}}_{j}|\right)\cdot p,$ (5)
where $Z$ is a standard Gaussian random variable and $p$ is a multiple testing correction factor. This approach sets the foundation for the baseline approach presented in Section 4.
Inference can also be made for the group of outcomes. This involves the following statistic, referred to as the Hotelling $T^{2}$ statistic, which generalizes the usual $t$-test:
@@
Under the null hypothesis $H_{0}:\tau=0$ that there is no treatment effect in any dimension, the Hotelling $T^{2}$ statistic in (6) follows the chi-squared distribution with $p$ degrees of freedom. Thus, the p-value for testing the null hypothesis $H_{0}$ is defined as
@@
where $\chi^{2}(p)$ is the chi-squared random variable with $p$ degrees of freedom. This approach is leveraged by our proposed approach of Section 5.
@@
As discussed in the introduction, we are primarily interested in settings where the treatment effect may be sparse within the $p$-dimensional outcome. Formally, throughout the remainder of the paper, we suppose that there exists a set of dimensions $S_{\tau}$ such that the sparsity index of $S_{\tau}$ is $|S_{\tau}|:=s_{\tau}<p$ and
$\mathbb{E}[Y_{j}|T=1,X]=\mathbb{E}[Y_{j}|T=0,X]\quad$ for $\text{for }j\notin S_{\tau},$ (8)
$\mathbb{E}[Y_{j}|T=1,X]\neq\mathbb{E}[Y_{j}|T=0,X]\quad$ for $\text{for }j\in S_{\tau}.$ (9)
Our goal in this paper is to present a practical regression-based method for approximate recovery of $S_{\tau}$, together with valid inference on treatment effects localized to the identified subset. Next, we present a simple, commonly employed baseline approach for this purpose (first without, then with pre-treatment covariates); we subsequently compare our proposed approach to this baseline.
@@ 4 A Baseline Approach
In this section, we present a simple, commonly used baseline approach for approximate recovery of $S_{\tau}$, that serves as a reference point.
4.1 Subset selection
A data scientist may be inclined to perform the dimension-by-dimension testing, with multiple testing correction, of Section 3 and then select all of the dimensions with significant p-values. If this approach is used with a pre-defined sparsity index (or bound on the sparsity index) $s_{\tau}$, then it is equivalent to ranking outcomes by their treatment effect sizes, and selecting outcome dimensions with the largest effect sizes. This provides the basis of a commonly used baseline approach and is equivalent to considering the following objective:
@@ 4.1 Subset selection
where
$\hat{\beta}=\arg\min_{\beta}\Bigg{|}\Bigg{|}\left(\frac{{\hat{\tau}_{\text{DiM}}}_{j}}{{\sqrt{\hat{\Sigma}_{\text{DiM}}}_{jj}}}\right)_{j=1}^{p}-\beta\text{}\Bigg{|}\Bigg{|}_{2}^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau}.$ (10)
Here, $||\beta||_{0}=\sum_{j=1}^{p}1\{\beta_{j}\neq 0\}$ is the $\ell_{0}$ norm of $\beta$. That is, $\hat{S}_{\tau}$ is simply the top $s_{\tau}$ outcomes based on the effect sizes of average treatment effects.
4.2 Inference via sample-splitting
We are not only interested in identifying the dimensions along which there may exist a treatment effect, but also performing inference along the selected dimensions. To accomplish this, we consider two sets of data: one for selection and one for inference. As discussed in the introduction, our data may naturally be divided in clinical settings by a POC study on which we select dimensions and a larger clinical trial on which we estimate treatment effects. Alternatively, we can deploy a sample-splitting approach on a single pool of data. Regardless, the above selection step is applied to the first data split.
Having identified a subset $\hat{S}_{\tau}$ on the first data split, inference can proceed as in Section 3 on the second data split, but restricted to the set $\hat{S}_{\tau}$. We can specialize the definitions in (1) and (4) to the subset $\hat{S}_{\tau}$, and carry out inference using the second data split. Although subset selection on the first split in this baseline approach was carried out using an approach inspired by (5), inference on the second split can proceed according to either (5) or (7) defined with these $\hat{S}_{\tau}$-specific quantities. Note that in (5), we would still apply a correction factor for multiple testing, but it is significantly smaller, especially when $s_{\tau}=|\hat{S}_{\tau}|$ is much smaller than $p$.
4.3 Incorporation of pre-treatment covariates
In causal inference, pre-treatment covariates are often leveraged to reduce the variance of the treatment effect estimators (Deng et al. [2013], Lin [2013], Jin and Ba [2023]). As an example, the CUPED estimator [Deng et al., 2013] estimates $\tau$ as